# PreAlign: Boosting Cross-Lingual Transfer by Early Establishment of Multilingual Alignment

### v2

@@ 3.1 Injecting Multilingual Alignment before Language Pretraining
$\mathcal{L}^{l}_{\text{align}}=\sum_{\begin{subarray}{c}w_{j}\in\mathcal{W}\\ w_{i}\in T(w_{j})\end{subarray}}\log\frac{\exp(cos(h^{l}_{w_{i}},h^{l}_{w_{j}})/\tau)}{\sum_{w_{k}\in\mathcal{B}}\exp(cos(h^{l}_{w_{j}},h^{l}_{w_{k}})/\tau)}$ (2)
@@ 5.3 Ablation Study
Joint Training 25.7 99.7 43.5 46.9 49.8 80.6 64.6 63.5 58.3 62.0 - 26.2 25.1 26.8 26.3
PreAlign 25.4 91.1 39.8 40.7 44.6 80.6 69.2 67.5 60.8 65.1 - 45.7 48.2 43.4 46.0
@@
Joint Training 20.3 79.8 32.5 34.8 39.6 82.3 65.8 65.3 56.9 63.7 - 37.8 39.5 36.1 37.7
PreAlign 19.9 75.2 28.3 30.7 33.6 82.4 70.0 69.3 65.6 68.2 - 63.8 66.5 64.7 63.6
@@
Joint Training 15.8 62.2 24.0 27.7 31.2 84.3 70.8 70.6 63.7 68.6 - 49.6 44.1 45.5 48.0
PreAlign 16.1 58.0 23.3 25.3 29.4 83.9 74.0 72.9 68.2 71.4 - 71.1 73.9 72.7 72.5

### v3

@@ 3.1 Injecting Multilingual Alignment before Language Pretraining
$\mathcal{L}^{l}_{\text{align}}=\sum_{\begin{subarray}{c}w_{j}\in\mathcal{W}\\ w_{i}\in T(w_{j})\end{subarray}}\text{log}\frac{\exp(\ d(h^{l}_{w_{i}},h^{l}_{w_{j}})/\tau)}{\sum_{w_{k}\in\mathcal{B}}\exp(\ d(h^{l}_{w_{j}},h^{l}_{w_{k}})/\tau)}$ (2)
@@ 5.3 Ablation Study
Joint Training 25.7 99.7 43.5 46.9 49.8 80.6 64.6 63.5 58.3 62.0 26.2 25.1 26.8 26.3
PreAlign 25.4 91.1 39.8 40.7 44.6 80.6 69.2 67.5 60.8 65.1 53.1 57.2 51.6 55.5
@@
Joint Training 20.3 79.8 32.5 34.8 39.6 82.3 65.8 65.3 56.9 63.7 37.8 39.5 36.1 37.7
PreAlign 19.9 75.2 28.3 30.7 33.6 82.4 70.0 69.3 65.6 68.2 63.8 66.5 64.7 63.6
@@
Joint Training 15.8 62.2 24.0 27.7 31.2 84.3 70.8 70.6 63.7 68.6 49.6 44.1 45.5 48.0
PreAlign 16.1 58.0 23.3 25.3 29.4 83.9 74.0 72.9 68.2 71.4 71.1 73.9 72.7 72.5